Can you help me solve this problem please!
topic:quadratic equations by completing the square
marivic is five years older than michelle. in three years , the prdocut of her age and michelle's age five years ago will be 90 years.Find their present ages.

Answer by mananth(16949)

(x-5)(x+8)=90
x^2+3x-40=90
x^2+3x=130
x^2+3x+(9/4)=130
(x+3/2)^2= 130
(x+3/2)= sqrt130
x+1.5 = +/- 11.4
x=-12.9 ( ignore)
OR
x=9.9 =10 Michelle's age
Marivic's age = 15
...
m.ananth@hotmail.ca